How many faces and edges does a triangular prism have?

Grade:12

3 Answers

Vikas TU
3 years ago
Dera student 
Traingular prism have 5 No of faces and total of 9 edges. 
Hope this will help 
Good Luck 
Cheers 
ankit singh
3 years ago
5 Faces
6 Vertices
9 Edges
Because it is a prism there are two identical bases, in this case triangles, and additionally there is one rectangular face for each side of the base figure, connection the two identical bases. For a prism the number of faces is always two more than the number of sides of the base (b + 2). (For a pyramid the number of faces is number of sides on the base + one)
So there are 5 faces, 3 + 2 = 5.
For vertices there is a 3 vertices on one base triangle and 3 more on the other base triangle. For a prism the number of vertices is always twice as many vertices as sides of the base (b + b or 2b). (For a pyramid it is base + one)
So there are 6 vertices, 3 + 3 = 6.
For edges there are 3 edges on the bottom triangle, 3 edges on the top triangle and 3 edges connecting the top with the bottom. For a prism the number of edges is always three times as many sides as the base (b + b + b or 3b). (For a pyramid it is base X two)
So there are 9 edges, 3 + 3 + 3 = 9.
ABD
2 years ago
Traingular prism have 5 No of faces and total of 9 edges. 
 
Hope it will help you.
Thanks 👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍👍